what is the conservation of 1.2 g of NaNO3 dissolved in 80.0 mL of solution


1
Expert's answer
2020-10-27T12:40:35-0400

In this context, I think what you wanted to ask is what is the concentration of 1.2 g of NaNO3 dissolved in 80.0 mL of solution

A quick way to calculate concentration is through this formula

Concentration = (mass/molar mass) ×(1000/volume in mL or cm3)

Molar mass of NaNO3=[23+14+(16×3)]=23+14+48=85g/mol

Concentration= (1.2/85)×(1000/80)= 0.176M


The concentration of the solution is 0.176M